OpenStudy (anonymous):

The net price equivalent rate of 9/15/18 is: a. .36573 b. .63573 c. .36427 d. .63427 e. None of these

jhonyy9 (jhonyy9):

(3/5)x(1/18)=3/90=1/30=,033 from this resulted e. none of these